### v1

# Advancing TTP Analysis: Harnessing the Power of Encoder-Only and Decoder-Only Language Models with Retrieval Augmented Generation

Abstract
Tactics, Techniques, and Procedures (TTPs) outline the methods attackers use to exploit vulnerabilities. The interpretation of TTPs in the MITRE ATT&CK framework can be challenging for cybersecurity practitioners due to presumed expertise, complex dependencies, and inherent ambiguity. Meanwhile, advancements with Large Language Models (LLMs) have led to recent surge in studies exploring its uses in cybersecurity operations. This leads us to question how well encoder-only (e.g., RoBERTa) and decoder-only (e.g., GPT-3.5) LLMs can comprehend and summarize TTPs to inform analysts of the intended purposes (i.e., tactics) of a cyberattack procedure. The state-of-the-art LLMs have shown to be prone to hallucination by providing inaccurate information, which is problematic in critical domains like cybersecurity. Therefore, we propose the use of Retrieval Augmented Generation (RAG) techniques to extract relevant contexts for each cyberattack procedure for decoder-only LLMs (without fine-tuning). We further contrast such approach against supervised fine-tuning (SFT) of encoder-only LLMs. Our results reveal that both the direct-use of decoder-only LLMs (i.e., its pre-trained knowledge) and the SFT of encoder-only LLMs offer inaccurate interpretation of cyberattack procedures. Significant improvements are shown when RAG is used for decoder-only LLMs, particularly when directly relevant context is found. This study further sheds insights on the limitations and capabilities of using RAG for LLMs in interpreting TTPs.
Introduction
The evolving changes of cybersecurity vulnerabilities presents a significant challenge for incident and threat analysis. Cyber Threat Intelligence (CTI) reports focus on gathering, analyzing, and interpreting data about potential cyber attacks. CTI aims to understand and anticipate the Tactics, Techniques, and Procedures (TTPs) of cyber adversaries. The process of deciphering and understanding the various cyberattack procedures (which is constantly updated in reference models like the MITRE ATT&CK [^1] https://attack.mitre.org/), requires extensive expertise and effort. The ATT&CK framework serves as a crucial tool in this regard, offering in-depth insights into adversary behaviors and techniques in targeting systems. The ATT&CK framework utilizes a 3/4-tier Tactic-Technique/Sub-technique-Procedure hierarchy to explain the how and why behind attackers’ exploitation of vulnerabilities. However, the complexity and potential ambiguity of these descriptions can make them challenging to interpret, leading to differing conclusions among security analysts from the same description. For example, consider the following attack procedure description:
Threat Group-3390 has performed DLL search order hijacking to execute their payload.
The ATT&CK framework associates three tactics with this procedure, namely, ‘Privilege-Escalation’, ‘Defense-Evasion’, and ‘Persistence’. It is not obvious that the DLL Search Order Hijacking technique will lead to any of the three, not to mention all three adversarial intents. In fact, some might argue that this description is more aligned with the ‘Execution’ tactic, as the adversary tries to execute its malicious payload. Therefore, we ask: Can LLMs interpret and map procedure descriptions to specific ATT&CK tactics? Note that we recognize the inherent ambiguity of ATT&CK TTP mapping. This in fact motivated us to study the capabilities and limitations of LLMs under such real-world imperfections.
The evolution of Large Language Models (LLMs), such as the OpenAI’s GPT-3.5 model [^2] https://platform.openai.com/docs/models, has led to substantial progress in NLP tasks (Min et al. 2023; Zhao et al. 2023). This progress is mainly due to the LLM’s enhanced semantic understanding and scalability. There have also been works that process TTPs by leveraging LLMs (Sauerwein and Pfohl 2022; Husari et al. 2017; You et al. 2022; Rani et al. 2023; Alves, Geraldo Filho, and Gonçalves 2022; Fayyazi and Yang 2023). However, none of these works compared the use of encoder-only LLMs (e.g., RoBERTa) and decoder-only LLMs (e.g., GPT-3.5) while leveraging Retrieval Augmented Generation (RAG) techniques to analyze the descriptions of attack procedures.
The encoder-only LLMs excel in tasks like sequence classification and entity recognition by processing input data to generate contextual representations, while decoder-only models adept at generating coherent, contextually relevant text, making them ideal for language generation tasks. One major problem of LLMs is the tendency to hallucinate. This issue involves the model’s ability to produce responses that are misleading or completely fabricated, yet presented confidently. To address this issue, RAG techniques are proposed to extract the most relevant documents and feed those as additional information into the input of LLMs. It is used to shift the model’s attention to relevant information by first retrieving documents that are likely to contain the answer and then using this information to generate a response. This method is beneficial as it is not only limited to model’s temporal knowledge but also plays a crucial role in reducing hallucinations by grounding responses from selected resources.
Recognizing the potentially different uses and benefits for encoder-only and decoder-only LLMs for cybersecurity operations, this study sets out to compare the two in their effectiveness to interpret cyberattack procedures and map them to ATT&CK tactics. Specifically, we examine three sets of approaches: 1) direct use of the decoder-only LLMs (GPT-3.5-Turbo), 2) supervised fine-tuning (SFT) of encoder-only LLMs (RoBERTa (Liu et al. 2019) and SecureBERT (Aghaei et al. 2022)) with ATT&CK descriptions, and 3) the use of RAG on decoder-only LLMs. Note that in all three approaches, we avoid explicitly training of LLMs to map the procedure descriptions to MITRE ATT&CK tactics. This is meant to test LLM’s interpretation capability instead of building a traditional classifier model. In a way, we hope to mimic, to the extent possible, how LLMs perform if the procedure description does not exist on the Internet.
To the best of our knowledge, this is the first formal study to compare these approaches to interpret cyberattack procedure descriptions. By experimenting using the latest TTP descriptions extracted from MITRE ATT&CK (release v14.1 as of Oct. 31, 2023), we present the benefits and limitations of the two LLM classes and suggest future research directions. Our main contributions are as follows:
  • We analyze and compare the performance in SFT of smaller encoder-only LLMs vs. RAG-enhanced of larger decoder-only LLMs for interpreting TTPs.
  • We demonstrate a significant improvement in interpreting TTPs for decoder-only LLMs when the right pieces of information (via RAG) is provided.
  • We study the different retrieval methodologies to provide relevant context of procedure descriptions and show the limitations of RAG techniques when they fail to retrieve relevant information.
  • We demonstrate that while decoder-only LLMs maintain strong ’recall’, they lack ’precision’, and we elaborate the need to new advances in increasing the ‘precision’ in LLMs without compromising ‘recall’.
  • We demonstrate the current capabilities of using RAG with decoder-only LLMs by providing specific examples, and offer insights into their lack of ‘interpretability’.
Related Works
Large Language Models
The advancements in Large Language Models (LLMs) are largely attributed to the adoption of the Transformer architecture (Vaswani et al. 2017). Central to this architecture are the encoder and decoder components, which have been utilized distinctively in various LLMs.
Encoder-Only: One of the pioneering introductions is BERT (Devlin et al. 2019), a pre-trained encoder-only LLM based on transformers. It leverages the Transformer encoder to achieve a deep understanding of language context and nuances. BERT’s significance lies in its bidirectional training, which enables a comprehensive understanding of language structure. Moreover, RoBERTa (Liu et al. 2019), extended BERT’s capabilities. It optimized hyperparameters, eliminated the Next Sentence Prediction task, and enhanced training with larger mini-batches and longer sequences. Another notable encoder-only model is SecureBERT (Aghaei et al. 2022), which is based on RoBERTa but specifically fine-tuned with cybersecurity data.
Decoder-Only Models: The landscape of LLMs has been further transformed by decoder-only models, characterized by their extensive training datasets and significant computational resources. OpenAI’s GPT-3.5 and Meta’s LLAMA-2333https://ai.meta.com/llama are leading examples in this category, both encompassing billions of parameters. These models have not only shown the ability to scale but also their proficiency in generating human-like text and recognizing intricate patterns across various fields. The effectiveness and versatility of these models in handling diverse and complex tasks have been extensively documented in the literature, as described in (Min et al. 2023; Zhao et al. 2023). However, there remains a question on how capable these models are in interpreting TTPs.
LLM for TTP interpretations
In the context of processing TTP descriptions, several works have been proposed by using NLP techniques. (Sauerwein and Pfohl 2022; Husari et al. 2017; Kim, Kim et al. 2022).
More recently, the following works have utilized encoder-only LLMs (e.g., BERT) for classifying TTPs. You et al. (2022) introduced the TIM framework, using continuous description sets and SentenceBERT embeddings, along with TTP-specific features like IPs and URLs, to categorize TTPs based on five techniques and one tactic of the ATT&CK framework. Alves, Filho, and Gonçalves (2022) tested various BERT model variants with different hyperparameters to find the the best model for TTP classification. Rani et al. (2023) developed TTPHunter, a tool that uses BERT and RoBERTa models with a linear classifier for extracting TTPs from unstructured text. Orbinato et al. (2022) conducted an experimental study to compare traditional ML and DL models (e.g., SVM vs. SecureBERT) for mapping TTPs from Cyber Threat Intelligence texts. The authors found SecureBERT to be the most effective, however, they did not provide a comparison between SecureBERT and its foundational RoBERTa model.
Furthermore, our previous work (Fayyazi and Yang 2023) compared the encoder-only vs the direct use of decoder-only LLMs to process and classify higher-level tactic and technique descriptions. The encoder-only experiments focused on traditional classifier training where training and testing sets are drawn from the same set of data. The decoder-only approach did not integrate with Retrieval Augmented Generation. The current work significantly extend our previous findings to treat harder-to-interpret attack procedure descriptions and with more advanced LLM approaches without explicitly mapping the testing data to tactic labels.
Retrieval Augmented Generation (RAG)
Despite the strong capabilities of state-of-the-art LLMs in instruction-following tasks (e.g., question-answering), these models have shown to be prone to hallucination in their generated responses. Their reliance on pre-trained knowledge (i.e., constrained by the training data’s temporal scope) poses limitations, particularly in domains where accuracy is paramount like cybersecurity. Fine-tuning with fresh data can be expensive and still prone to hallucination. Therefore, to adapt to new information and to provide factual knowledge, RAG was introduced (Borgeaud et al. 2022).
In RAG, relevant pieces of information is stored in a vector database. When a query is presented, RAG selects the most relevant data from this database, which is then incorporated into the prompt of LLMs before they generate a response. This approach offers two key advantages: first, it allows the LLM to incorporate current information without needing constant fine-tuning; second, it enhances the clarity and relevance of the responses by providing contextually grounded information. Essentially, RAG equips LLMs with the ability to gain a focused understanding of the specific context of a query. This approach improves the clarity and precision of the responses by directing the model’s attention towards the most relevant context retrieved. There have been works that used RAG in NLP tasks (Borgeaud et al. 2022; Mallen et al. 2023; Ram et al. 2023; Al Ghadban et al. 2023). In this work, we propose the use of RAG to provide more information about an attack procedure to the LLM and focus the LLM’s attention towards the text retrieved through RAG.
Methodology & Experimental Design
We design this study to compare: 1) supervised fine-tuning (SFT) of encoder-only LLMs with labeled ATT&CK technique/sub-technique descriptions, 2) direct use of decoder-only LLMs (i.e., pre-trained knowledge), and 3) use of RAG for decoder-only LLMs with relevant URLs retrieved by finding the most similar attack procedures. Furthermore, we consider an ideal (yet unrealistic) case where the exact URL of each procedure is used for RAG to give the upper bound performance. This experimental design enables us to assess how well LLMs interpret cyberattack procedure descriptions and map them to the corresponding tactic(s). This is a challenging process due to multiple reasons: 1) there are no publicly available datasets that map procedure descriptions into ATT&CK tactic(s), 2) each procedure can map to multiple ATT&CK tactics that LLMs need to consider, 3) a well-structured prompt is needed for LLMs to generate desired outputs, and 4) RAG needs to retrieve the most relevant information. We discuss next our methodology and design of experiments to tackle these challenges.
Figure 1: The process of supervised multi-label fine-tuning of encoder-only LLMs with high-level ATT&CK descriptions (excluding ‘procedures’) – adopted from (Fayyazi and Yang 2023).
Datasets
To conduct our experiments, we gathered data from the MITRE ATT&CK framework. We chose ATT&CK due to its broad adoption in industry SIEM tools and associated detailed cybersecurity descriptions. We curated the descriptions of enterprise tactics, techniques, and sub-techniques along with the mapping to their corresponding tactic(s) for fine-tuning the encoder-only models. In total, we obtained 639 descriptions from the MITRE ATT&CK framework. Some descriptions have 2,3, or even 4 different tactics. Note that the use of 639 labeled ATT&CK descriptions for supervised fine-tuning of encoder-only LLMs reflects the real-world scenario where only limited labeled data would exist. Moreover, to test the performance of both encoder-only and decoder-only LLMs, we crawled all the enterprise procedure examples that describe a attack methodology from the ATT&CK framework. We remove the procedure descriptions that contain any of the 14 MITRE ATT&CK tactics within the descriptions to prevent potential biases or shortcuts of the LLMs. In total, we gathered 9,532 procedure descriptions along with their corresponding URLs (for RAG). Note that the procedure descriptions are not used for SFT of encoder-only LLMs. Figure 2 shows the pair-wise overlap across tactics for attack procedure descriptions. As can be seen, ‘Persistence’, ‘Privilege-Escalation’ and ‘Defense-Evasion’ tactics have higher overlaps with the others. We will share the curated data along with our source code as artifacts for the research community.
Figure 2: The pair-wise overlap of tactics on entire ATT&CK procedure descriptions.
Supervised Fine-Tuning of Encoder-Only LLMs
To conduct supervised fine-tuning with encoder-only LLMs, we selected RoBERTa-base (Liu et al. 2019) and SecureBERT (Aghaei et al. 2022). We chose these models because of their demonstrated superiority over traditional ML and DL techniques in processing TTPs (Rani et al. 2023; Alves, Geraldo Filho, and Gonçalves 2022; Orbinato et al. 2022). The SecureBERT model has already fine-tuned using an extensive cybersecurity text corpus. Our goal is to assess these LLMs’ efficacy in distilling crucial information from cybersecurity descriptions and mapping them into relevant ATT&CK tactics. Therefore, we fine-tune them with the ATT&CK’s Tactics, Techniques, and Sub-techniques descriptions, but not procedures. It is worth noting that both RoBERTa and SecureBERT models have the limitation to 512-token input limit, therefore, we truncated the ends of the descriptions that exceeded this limit. Figure 1 shows the overall process of fine-tuning both RoBERTa and SecureBERT models. First, we tokenize the ATT&CK descriptions for compatibility with both RoBERTa and SecureBERT models. Second, we add a classification layer with 14 nodes with each corresponding to an ATT&CK tactic. To address the challenges of multi-label classification, we employ a Sigmoid activation function within this layer. This function is crucial as it calculates the independent probability of each tactic for a given description. This enables the models to effectively discern and assign multiple labels where appropriate.
Table 1: The hyperparameters tested and selected (bold numbers) for SFT of RoBERTa and SecureBERT.
Hyperparameter Value
Optimizer AdamW
Loss Func. BCE
Activation Func. Sigmoid
Batch Size [8, 16]
Epoch [25, 30, 35]
Learning rate [1e-5, 2e-5, 5e-5, 7e-7]
Table 1 shows the hyperparameters selected to fine-tune the models. We employed binary Cross-entropy as the loss function (due to the binary nature to predict each tactic), batch size of 16, 30 epochs, and 5e-5 for learning_rate as the best-performing value. The models’ output takes the form of a 14-dimensional binary vector, each element representing an ATT&CK tactic, with a Sigmoid output over 0.5 indicating a prediction for that specific tactic.
Decoder-Only LLMs with and w/o RAG
We consider OpenAI’s GPT-3.5-turbo-1106 model [^4] https://platform.openai.com/docs/models, which has a context window of 16K token, as our decoder-only LLM. It is worth noting that in OpenAI models, their inherent design are non-deterministic. This means that identical inputs can result in different outputs. Therefore, to manage this aspect and have consistency, we have set the temperature parameter to “0” and a seed number (1106) to ensure reproducibility. Our baseline use of GPT-3.5-Turbo will be without RAG. We engineered a simple and clean prompt without overly influencing the attention of the LLM to the prompt other than the procedure description, while attempting to avoid hallucination based on best practices reported in the community. Below is the engineered prompt used for the decoder-only baseline:
You are a cybersecurity expert.
Knowing that <<much-less-than<<< <procedure>>much-greater-than>>> >, what MITRE ATT&CK tactics will a cyber adversary achieve with this technique?
Please only respond with the MITRE ATT&CK tactics you are certain about.
We design our decoder-only LLM with RAG with an attempt to mimic a real-world scenario where the procedure in question does not exist and cannot be searched directly to find the corresponding tactic(s). We utilize FAISS [^5] https://github.com/facebookresearch/faiss to find top-3 similar procedures in our dataset for each procedure tested. These 3 similar procedures will provide up to 3 URLs where RAG will use to retrieve chunks of text serving as contexts for the prompt. We elected to use 3 chunks of 8,000 characters with chunk_overlap of 500 (using the langchain [^6] https://www.langchain.com/ framework), and store them in a Vector Store. Then, we use the OpenAI embeddings for both the question and the stored chunks, and we retrieve the top-3 most relevant chunks based on the question and insert them as ‘Relevant Context’ into the prompt. The prompt template used for GPT-3.5-Turbo is as follows:
You are a cybersecurity expert. Consider the relevant context provided below and answer the question.
Relevant Context:
{context}
Question: Knowing that <<much-less-than<<< <procedure>>much-greater-than>>> >, what MITRE ATT&CK tactics will a cyber adversary achieve with this technique?
Please only respond with the MITRE ATT&CK tactics you are certain about.
Note that in this case, some of the similar procedures can have the same URL as the procedure in question. This could lead to GPT-3.5-Turbo finding direct answer through RAG. However, as we will demonstrate in the results, the performance is good but still not perfect even if the exact URL is among the ones found. We also note that the ATT&CK’s techniques/sub-techniques URLs actually contain the tactic names (see the screenshot shown in Figure 3).
Figure 3: Screenshot of an ATT&CK URL [^8] https://attack.mitre.org/techniques/T1484that contains the associated tactics of a particular technique/sub-technique.
In order to provide the ideal RAG performance as reference, we also tested an ideal (yet unrealistic) scenario where the exact URL where the procedure in question resides is the sole URL provided for RAG. This enforce the selection of top-3 8,000 character chunks to come from the most relevant URL. We note that some URLs have less text and others have many. This variation may also influence the performance of RAG. Nevertheless, as we will show in the results, this ideal case achieves excellent upper bound performance.
Figure 4: The overall process of RAG based on similar ATT&CK procedures of decoder-only LLMs.
Results & Discussion
We first evaluate which encoder-only model performs better. Then, we compare the decoder-only models with and without RAG. Next, we compare the capabilities and limitations of the selected encoder-only and decoder-only LLMs with RAG. Finally, we provide specific examples for insights.
This work considers the ‘Samples Average F1’ score. F1 score captures the harmonic mean of Precision and Recall. We consider ‘Recall’ to assess LLM’s capability to interpret procedures and map them to the presumed tactics, and ‘Precision’ to assess the extent of LLM’s hallucination. The Samples Average F1 Score calculates the F1 score for each instance and then averages these scores. The reason for using the Samples Average, as opposed to Micro, Macro and Weighted average scores, is that it lessen the biased due to imbalanced number of procedures across tactics or over-emphasis of procedures that have multiple labels.
Evaluation of SFT of Encoder-Only LLMs
We fine-tune the encoder-only LLMs (RoBERTa-SFT and SecureBERT-SFT) using the curated MITRE ATT&CK enterprise Tactics, Techniques, and Sub-techniques descriptions (mapped to AT&CK’s tactics). We should emphasize that we do not fine-tune these models with procedure descriptions, as we want to evaluate the ‘interpretation’ capability of these LLMs (i.e., by not giving these models the explicit procedure, tactic mappings). For the results, we obtained 0.54 and 0.41 as Samples Average F1 score for SecureBERT and RoBERTa, respectively. SecureBERT exhibited stronger performance compared to the RoBERTa model, and is chosen to compare to the decoder-only LLMs.
Prompt-Only vs. RAG of GPT-3.5-Turbo
Table 2 shows the performance of the aforementioned test cases for all 9,532 ATT&CK procedure descriptions separated by tactics they are associated with according to ATT&CK (recall Fig. 2). Note that since some procedures are associated with multiple tactics, the sum of all supports (10,952) exceeds the number of tested procedures (9,532). As the results suggest, using only the prompt (i.e., relying on GPT-3.5’s pre-trained knowledge) results in a poor performance in interpreting the ATT&CK tactics with ‘0.60’ Samples Avg. F1 (but higher than SecureBERT-SFT). In 4 out of 14 tactics, it achieved a F1 score higher than ‘0.50’, but got ‘0.00’ for Resource Development. In contrast, when using RAG by providing the exact URL (the ideal case), the model’s performance for Samples Avg. F1 significantly improved to ‘0.95’, with 11 out of 14 tactics got F1 higher than ‘0.90’. This indicates that by providing the model with the accurate information (chunks of data that contain the answer), it achieves significantly better performance without the need for fine-tuning. However, in a more realistic scenario (RAG with top-3 similar procedure), where the exact piece of information of a ‘new’ attack procedure are not easily accessible, the performance, while still superior than prompt-only, falls down to ‘0.68’ Samples Avg. F1 Score. This implies that one should have limited expectation for LLM to interpret new, likely unseen attack procedures even with RAG .
[…]
Table 2: Per-tactic performance of Prompt-Only vs. RAG w/Exact URL vs. RAG w/ Similar Procedures with the GPT-3.5-Turbo model for all ATT&CK descriptions.
GPT-3.5-Turbo   Prompt-Only   RAG w/ Reference URL   RAG w/ Relevant Procedures
Tactics   F1 Score   F1 Score   F1 Score   Support
Collection   0.44   0.84   0.53   852
Command and Control   0.52   0.96   0.59   705
Credential Access   0.59   0.92   0.69   615
Defense Evasion   0.66   0.96   0.76   2669
Discovery   0.86   0.98   0.85   2342
Execution   0.48   0.90   0.55   1208
Exfiltration   0.20   0.65   0.31   92
Impact   0.44   0.84   0.54   213
Initial Access   0.42   0.91   0.46   406
Lateral Movement   0.45   0.90   0.44   212
Persistence   0.40   0.92   0.49   549
Privilege Escalation   0.41   0.92   0.60   723
Reconnaissance   0.21   0.99   0.42   91
Resource Development   0.00   0.93   0.50   275
Samples Avg. F1   0.60   0.95   0.68   10952
Decoder-Only w/ RAG vs. SFT of Encoder-Only
We further compare the SFT of Encoder-only (SecureBERT) vs. Decoder-only with RAG models by analyzing their performance based on Precision, Recall, and F1 scores. For Decoder-only with RAG, we separate the 9,532 procedures into two sub-groups: 1) at least one top-3 URL match to the exact URL of the procedure in question, and 2) no top-3 URL matches to the exact URL. This separation gives a detailed insights on how RAG performs with and without the exact URL, and to provide a comparison on the ‘main’ differences between these encoder-only and decoder-only models. Table 3 shows their performances. The encoder-only models perform a lot better in ‘Precision’ than ‘Recall’. This means that SecureBERT (the encoder-only LLM) tries to provide precise responses. However, as shown in GPT-3.5 (the decoder-only model) results, they perform significantly better for ‘Recall’ than ‘Precision’. This indicates that hallucinations in LLMs typically arise during the decoding phase, where the model tries to generate information not strictly present/implied in the input. This is a major factor that can reduce precision. Therefore, this indicates a need for improving ‘Precision’ in LLMs without compromising ‘Recall’. In addition, by comparing Prompt-Only results (table 2) and w/o Matched URLs (table 3), we can see when the LLM is provided with additional ‘distracting’ context (RAG), it is performing worse than relying solely on the model’s trained knowledge. This indicates that RAG puts high attention to the retrieved context (which is inserted into its prompt) and does not pay attention to the ‘correct’ answer (as it tries to look for the best suitable chunk of the additional context).
Overall, the results shown in Table 2 and Table 3 indicate that decoder-only LLMs with RAG (GPT-3.5) generally outperform the encoder-only LLMs (SecureBERT). However, it is important to note that encoder-only LLMs require significantly fewer resources for training and evaluation compared to their decoder-only counterparts. This distinction is crucial for understanding the trade-offs in efficiency and effectiveness between these two model architectures.
Table 3: Per-tactic performance of Encoder-Only and Decoder-Only for ATT&CK procedure descriptions. For GPT-3.5, we compare cases where the source and retrieved procedures share the same URL (4115 of 9532 procedures) and where they do not (5417 of 9532 procedure). Note that these numbers are higher for the tactics due to the multi-label situation.
Model:   GPT-3.5 SecureBERT
Data:   w/ Matched URLs w/o Matched URLs SFT w/ Tactic & Techniques
Tactics   Precision Recall F1 Score Support   Precision Recall F1 Score Support   Precision Recall F1 Score Support
Collection   0.58 0.96 0.72 354   0.31 0.67 0.42 498   0.56 0.56 0.56 852
Command and Control   0.73 0.96 0.83 314   0.33 0.55 0.41 391   0.74 0.26 0.39 705
Credential Access   0.79 0.99 0.88 268   0.46 0.71 0.56 347   0.69 0.71 0.70 615
Defense Evasion   0.87 0.97 0.92 1162   0.59 0.71 0.65 1507   0.80 0.60 0.68 2669
Discovery   0.95 0.97 0.96 1111   0.73 0.76 0.74 1231   0.95 0.69 0.80 2342
Execution   0.71 0.89 0.79 548   0.31 0.54 0.39 660   0.63 0.48 0.54 1208
Exfiltration   0.30 0.96 0.46 28   0.16 0.67 0.26 64   0.41 0.42 0.42 92
Impact   0.66 0.96 0.78 118   0.23 0.59 0.33 95   0.83 0.23 0.35 213
Initial Access   0.57 0.96 0.71 156   0.24 0.59 0.34 250   0.62 0.27 0.37 406
Lateral Movement   0.51 0.97 0.67 71   0.24 0.48 0.32 141   0.46 0.35 0.40 212
Persistence   0.70 0.93 0.80 172   0.30 0.49 0.37 377   0.43 0.42 0.42 549
Privilege Escalation   0.89 0.82 0.86 261   0.49 0.41 0.45 462   0.65 0.34 0.44 723
Reconnaissance   0.50 1.00 0.67 2   0.41 0.42 0.41 89   1.00 0.04 0.08 91
Resource Development   0.75 0.66 0.70 70   0.48 0.38 0.42 205   0.85 0.17 0.28 275
Samples Avg. F1   0.85 0.95 0.88 4635   0.47 0.65 0.52 6317   0.55 0.54 0.54 10952
Specific Examples for Decoder-only w/ RAG
The application of LLMs to the interpretation of TTPs poses several challenges. In this part, we show how the decoder-only LLMs deal with the retrieved texts.
Example 1 Procedure in question (tactic: Credential Access): Lslsass can dump active logon session password hashes from the lsass process. Top-3 similar procedures: 1. Mafalda can dump password hashes from LSASS.exe. 2. Earth Lusca has used ProcDump to obtain the hashes of credentials by dumping the memory of the LSASS process. 3. CrackMapExec can dump hashed passwords from LSA secrets for the targeted system. Retrieved context: …Adversaries attempt to access credential material stored in the process memory of the Local Security Authority Subsystem Service (LSASS) …Tactic: Credential Access…
In this example, the URL of the ‘procedure in question’ is one of the URLs associated with the top-3 similar procedures. Out of these 3 retrieved URLs, all 3 were describing a technique/sub-technique of Credential Access, which matches to the tactic the procedure in question is associated with. Such a scenario exemplifies an ideal situation where RAG successfully provided the decoder-only LLM (GPT-3.5) with the correct contextual information.
Example 2 Procedure in question (tactic: Execution): HOPLIGHT has used svchost.exe to execute a malicious DLL. Top-3 similar procedures: 1. Hydraq uses svchost.exe to execute a malicious DLL included in a new service group. 2. Higaisa named a shellcode loader binary svchast.exe to spoof the legitimate svchost.exe. 3. Tropic Trooper has injected a DLL backdoor into dllhost.exe and svchost.exe. Retrieved Context: …Execution via DLL injection may also evade detection from security products since the execution is masked …Tactics: Defense Evasion, Privilege Escalation…
For this example, the URL of the ‘procedure in question’ also matched within the URLs of top-3 similar procedures. However, this time the retrieved chunks includes technique/sub-technique descriptions of ‘Execution’, ‘Defense-Evasion’, and ‘Privilege-Escalation’. GPT-3.5 incorrectly mapped this procedure to ‘Defense-Evasion’ and ‘Privilege-Escalation’ tactics but not ‘Execution’. We noticed that the following information were present in one of the chunks: Tactics: Defense Evasion, Privilege Escalation. We believe that the model paid too much attention on these specific keywords, leading to incorrect tactical mappings. Interestingly, this occurred despite the fact that some sentences discussed about ‘Execution’ via DLL. This example shows GPT-3.5’s lack of true comprehension capability of provided context, as it seems that the model over-prioritize certain keywords and overlooks the broader context of the content.
Example 3 Procedure in question (tactic: Defense Evasion): Chimera has cleared event logs on compromised hosts. Top-3 similar procedures: 1. Chimera has staged stolen data locally on compromised hosts. 2. Chimera has used compromised domain accounts to gain access to the target environment. 3. Chimera has copied tools between compromised hosts using SMB. Retrieved Context: …Monitor processes that appear to be reading files from disparate locations and writing them to the same directory or file may be an indication of data being staged …
In this example, the URL of the ‘procedure in question’ did NOT match to any URLs of the top-3 similar procedures. GPT-3.5 correctly predicted this procedure as ‘Defense-Evasion’ even though it did not have access to the reference URL where the source procedure was originally located. In the retrieved context for this procedure, there were no mention to the ‘Tactics:’ keyword (unlike examples 1 and 2), and no exact mention to the 14 ATT&CK tactics. However, the GPT-3.5 model interpreted this procedure correctly based on the retrieved context (as shown in the example). This indicates some interpretation capability by the model, even when there are no explicit textual cues or predefined categories available.
Example 4 Procedure in question (tactic: Resource Development): GALLIUM has used a variety of widely-available tools, which in some cases they modified to add functionality and/or subvert antimalware solutions. Top-3 similar procedures: 1. GALLIUM has used stolen certificates to sign its tools including those from Whizzimo LLC. 2. GALLIUM leveraged valid accounts to maintain access to a victim network. 3. GALLIUM used a modified version of HTRAN in which they obfuscated strings such as debug messages in an apparent attempt to evade detection. Retrieved Context: …TA416 Goes to Ground and Returns with a Golang PlugX Malware Loader …
For this example, the URL of the ‘procedure in question’ also did NOT match to any URLs of the top-3. The retrieved context described a technique/sub-technique of ‘Resource-Development’ tactic. GPT-3.5 completely hallucinated for this procedure by predicting 10 out of 14 ATT&CK tactics (without including Resource Development). In the retrieved chunks for this procedure, there were no mention to the keywords ‘Tactic:’ and ‘Resource Development’. We also noticed that the retrieved information were not useful regarding the ‘procedure in question’. This shows that GPT-3.5 model failed to reason for interpreting the procedure description (when provided less relevant context).
[…]
Concluding Remarks
We conducted a comparative analysis of the performance in supervised fine-tuning of smaller encoder-only LLMs against RAG-enhanced larger decoder-only LLMs. This comparison focuses on determining which model type is more effective in analyzing cyberattack procedure descriptions. We observed a significant improvement in the interpretation of TTPs by decoder-only LLMs when supplemented with appropriate RAG inputs. This demonstrates the impact of the right information on LLM’s performance.
Furthermore, our research reveals the varied methodologies of RAG in refining TTP descriptions, while demonstrating the limitations of RAG techniques, particularly when they fail to retrieve the most relevant information. We indicated that current decoder-only LLMs, while exhibiting high recall, often lack precision. This observation underscores the need for new advancements in enhancing the precision of LLMs without sacrificing their recall capabilities. Moreover, we also demonstrated the lack of‘interpretation’ of decoder-only LLMs when provided with relevant context by providing analysis with the specific examples. In this case, we conjecture that LLMs attempt to provide responses by searching for what they consider to be the most probable solution.